# A Survey of Large Language Model Empowered Agents for Recommendation and Search: Towards Next-Generation Information Retrieval

### v1

@@ Abstract.
Information technology has profoundly altered the way humans interact with information. The vast amount of content created, shared, and disseminated online has made it increasingly difficult to access relevant information. Over the past two decades, search and recommendation systems (collectively referred to as information retrieval systems) have evolved significantly to address these challenges. Recent advances in large language models (LLMs) have demonstrated capabilities that surpass human performance in various language-related tasks and exhibit general understanding, reasoning, and decision-making abilities. This paper explores the transformative potential of large language model agents in enhancing search and recommendation systems. We discuss the motivations and roles of LLM agents, and establish a classification framework to elaborate on the existing research. We highlight the immense potential of LLM agents in addressing current challenges in search and recommendation, providing insights into future research directions. This paper is the first to systematically review and classify the research on LLM agents in these domains, offering a novel perspective on leveraging this advanced AI technology for information retrieval. To help understand the existing works, we list the existing papers on agent-based simulation with large language models at this link: https://github.com/tsinghua-fib-lab/LLM-Agent-for-Recommendation-and-Search.
@@ 1. Introduction
Information technology has fundamentally transformed human life, particularly in how we interact with information. The development of the information age has led to the creation, sharing, and dissemination of vast amounts of content on the internet, making it increasingly challenging for individuals to access relevant information. Against this backdrop, the past two decades have witnessed the evolution of information systems, including search systems (Brin and Page, 1998; Speretta and Gauch, 2005) and Recommender Systems (RSs) (Zhang et al., 2019; Gao et al., 2023b). Typically, search systems passively retrieve relevant content from vast amounts of information with users’ query words, while recommender systems actively generate a list of candidates that users might be interested in based on collected user behavioral data or profile information (Schütze et al., 2008; Aggarwal et al., 2016; Levene, 2011). Some researchers consider recommendation as a specialized form of search, where the understanding of potential user needs is modeled as search terms (Xu et al., 2018).
@@
Recently, large language models (LLMs) have achieved remarkable success and are considered one of the most viable paths toward general artificial intelligence (Achiam et al., 2023). Large language models have demonstrated capabilities similar to or even surpassing human performance, not just in language-related tasks, exhibiting general understanding, reasoning, and decision-making abilities (Zhao et al., 2023b). Researchers have leveraged the text processing and commonsense reasoning capabilities of large models to conduct a series of search and recommendation tasks (White, 2024; Zhu et al., 2023; Ai et al., 2023). Furthermore, to overcome the limitations of large language models, researchers build various large language model agents (Xi et al., 2023). Specifically, large language model agents use large language model as the core and are equipped with functions such as memory management, workflow, input-output interfaces, external tool invocation, and large-small model collaboration, making their intelligence levels closer to human and capable of handling more complex tasks (Zhao et al., 2024a; Park et al., 2023; Zhou et al., 2023c).
@@
For the crucial and fundamental tasks of search and recommendation, researchers increasingly recognize the high value of large language model agents. In this paper, we make the first attempt to provide a systematic and comprehensive review of the research efforts on search and recommendation with large language model agents. We first introduce relevant background knowledge and further discuss the motivations for utilizing LLM agents, as well as how LLM agents address those critical challenges faced by search and recommendation. Next, we establish a taxonomy, categorizing these recent advances according to the role of large language model agents in the overall search or recommender system. We then elaborate on these works and analyze how they construct and use LLM agents. It is worth mentioning there have already been several works discussing the application of LLMs in the fields of IR and search, which either focus solely on LLMs (Chen et al., 2024a; Bao et al., 2024; Wu et al., 2024b; Gao et al., 2024) or a specific IR application (Peng et al., 2025), but this survey paper significantly differs by focusing on the application of LLM agents across the entire spectrum of IR functionalities.
@@
  • First, this paper is the first to review and organize the research on large language model agents in the research of search and recommendation, which is a new yet fast-growing research field.
@@
The structure of this paper is as follows: In Section 2, we introduce the background knowledge of recommender systems, search systems, large language models, and large language model agents. In Section 3, we analyze why large language model agents are necessary for search and recommendation. In Section 4, we first establish a coherent and comprehensive taxonomy system and then elaborate on the existing work. In Section 6, we systematically analyze the current problems and significant future directions for search and recommendation based on large language model agents. Finally, we conclude the paper in Section 7.
@@ 2.1.1. Interaction Interface
The rapid development of smart devices and mobile Internet and has exposed users to numerous applications employing recommender systems and search engines. The traditional interaction interface between users and search engines usually consists of a keyword query submission from users (e.g., on a search engine website or a project management software) and a one-round result retrieval from the search engines (Baeza-Yates et al., 1999). Interacting with a recommender system is even simpler, where users are mostly unconsciously browsing contents (e.g., on an e-commerce website, or a video sharing platform) provided by the recommender system (Resnick and Varian, 1997). Conversational recommender systems and search engines have recently gained attention (Zhang et al., 2018). In this interface, users interact through a chat interface using natural language, iteratively providing additional information or feedback based on previous results, forming a multi-turn dialogue. This conversational interface is more closely aligned with human communication methods and achieves more precise and satisfying results.
@@ 2.1.2. User/Query Modeling
For recommender systems and search engines to retrieve satisfying results, they must understand the users’ intentions. For example, identifying the products a user wants to buy on an e-commerce app or determining relevant websites based on a keyword query. This requires precise modeling of user characteristics and input queries. In recommender systems, this includes utilizing the user profile (gender, age, preferences, etc.), browsing/clicking/purchasing history, and more (Lops et al., 2011). In search engines, this involves text segmentation, named entity recognition, part-of-speech tagging, relation extraction, etc (Croft et al., 2010). Advanced systems like multi-modal and conversational search engines also incorporate image recognition, multi-turn dialogue comprehension, and other algorithms (Jannach et al., 2021; Wang et al., 2016).
@@ 2.1.3. Item Modeling
Apart from user/query modeling, item modeling also plays an essential role in efficient and extensive retrieval. It gathers information of the items, and preprocesses them into some representation (e.g., a feature vector) for later usage. In traditional search engines, this process is known as Indexing (Croft et al., 2010), where significant keywords or tags from web page content are identified and stored beforehand for faster retrieval. Advanced search engines may also use auxiliary knowledge bases specific to certain domains. In recommender systems, item modeling considers item descriptions, attributes, images, buyers, and other related information (Lops et al., 2011; Liu et al., 2024).
@@ 2.2.1. Large Language Models
Language models play a pivotal role in Natural Language Processing (NLP), primarily used for assessing sentence probabilities and generating grammatically correct text. In its early stages, the development of language models relied on statistical and probability theories such as n-grams, maximum entropy, and Hidden Markov Models, but was limited by its ability to handle long-range dependencies, known as the ”curse of dimensionality.” With the rise of deep learning technologies, neural network language models gradually became mainstream due to their robust generalization capabilities and flexible architectural designs. Models like Recurrent Neural Networks (RNNs) (Sherstinsky, 2020), Long Short-Term Memory networks (LSTMs) (Graves and Graves, 2012), and Gated Recurrent Units (GRUs) (Cho et al., 2014) effectively mitigated the challenges of processing long sequences by utilizing internal memory units, thereby better capturing contextual dependencies within sequences and significantly enhancing language model performance.
The introduction of the Transformer model (Vaswani et al., 2017) revolutionized the NLP field by replacing recurrent structures with self-attention mechanisms, thereby improving training speed and parallelism. Concurrently inspired by the pre-training and fine-tuning paradigm from computer vision, language models began leveraging large amounts of unlabeled data for pre-training, followed by task-specific fine-tuning, which gave rise to high-performance models such as BERT (Devlin et al., 2018). With the rapid expansion of training data and significant growth in computational resources, LLMs have established absolute dominance in the NLP. This wave not only symbolizes the comprehensive transformation of NLP from focusing on single-task optimization to multimodal perception and generation but also heralds language models reaching new heights in generation, understanding, and transformation. In the latest research, LLMs such as GPT-4 (Achiam et al., 2023) Claude (Anthropic, 2024), etc. not only perform well in tasks across various NLP tasks such as text generation, multilingual translation, and sentiment analysis but also demonstrate impressive performance in tasks traditionally considered to require deep cognitive abilities like solving mathematical problems, logical reasoning, and strategic planning. This comprehensive performance marks the maturity of artificial intelligence technology, steadily advancing in a direction closer to Artificial General Intelligence (AGI).
@@ 2.2.2. LLM Agents
At the beginning of the 21st century, breakthroughs in machine learning, particularly represented by the rise of deep learning, enabled autonomous learning from massive amounts of data rather than relying solely on pre-set programming rules. This marked a shift in agent research from rule-based methods to data-driven approaches. Deep neural networks significantly enhance the capabilities of intelligent agents in areas such as image recognition, speech processing, and natural language understanding, thus ushering in a new era of Artificial Intelligence (AI) agents. Recently, DeepSeek (Guo et al., 2025) and GPT series (OpenAI, 2022; Achiam et al., 2023) and other major language models have made significant progress in NLP, greatly improving the performance of artificial intelligence agents and demonstrating their profound potential.
@@
  • Perception Module: The perception module is equivalent to the agent’s ”senses”, ensuring that the agent can obtain the information needed for decision-making on time. Unlike conventional single-channel information processing systems, LLM agents demonstrate superior capability in seamlessly integrating multimodal information sources, including textual, auditory, visual, and even video data, thereby establishing a more robust and comprehensive perceptual framework (Schumann et al., 2024).
  • Control Module: The control module plays the role of a proxy ”brain”, analyzing the collected information, evaluating possible action plans, and selecting the best strategy. In LLM agents, an LLM serves as the core of the control module, enhancing the agent’s understanding and generation ability of language contexts and strengthening its situational adaptability, creativity, and long-term memory functions.
  • Action Module: The action module is equivalent to the ”limbs” of the agent, ensuring that the agent’s decisions are implemented, thereby changing the environmental state or achieving specific goals. In the LLM agent, this module is not limited to simple execution actions. Still, it is endowed with deeper functions, which can realize complex tasks such as language generation, tool use, and even concrete actions, greatly expanding the application scope and influence of the agent.
@@
3. Why LLM agents can be used for search and recommendation
@@ 3.2. LLM Agents can Interact with Environment and Integrate Information
In the field of information retrieval, another significant advantage of LLM agents is their ability to interact with environments and integrate results (He et al., 2025; Zhou et al., 2023a). IR often involves multiple data sources and complex retrieval scenarios, where traditional retrieval systems may struggle to effectively collect information from different sources. LLM agents with action modules can use various tools, browse web pages, operate mobile apps, and even independently search the internet to find needed information (Schick et al., 2024; Qin et al., 2023). Additionally, LLM agents can go a step further by filtering and summarizing the collected content, presenting users with a clear and concise report instead of delivering raw, unprocessed results like traditional IR systems. For example, when conducting academic literature searches, an LLM agent can interact with multiple academic databases to obtain research outcomes from different fields. It then analyzes and consolidates these results to provide users with a comprehensive literature review. Furthermore, LLM agents can interact with users, continuously adjusting retrieval strategies based on user feedback, thereby enhancing the quality of the retrieval results.
@@ 3.3. LLM Agents Can Serve as User Simulators to Generate Feedback to Information System
IR systems rely on real user feedback for improvement, but trial and error with actual users is costly and can negatively impact user experience. The various modules within LLM agents enable them to simulate human perception, memory, and actions, making them well-suited for collecting feedback in place of real users. LLM agents can simulate real user responses, making them invaluable for testing user experience when using a product and identifying potential improvements (Zhang et al., 2024g). They can also be used to explore the evolution of users’ mindsets and interests, aiding researchers in better understanding user behavior (Park et al., 2023; Takata et al., 2024). LLM agents can simulate users’ preferences and interact with recommender and search systems. These generated interaction data can be used to train recommender and search models, which can lead to more robust models (Sekulić et al., 2024b; Zhang et al., 2024f).
In summary, LLM agents bring several critical capabilities to the table that make them highly suitable for search and recommendation tasks. Their ability to think and plan, interface with various tools, and simulate user behavior makes a significant difference in the user experience of IR in the new era.
@@ 4. Recent Advances of LLM Agents for Recommendation and Search
Recently, significant advancements have been made in the research on combining LLM agents with traditional search and recommendation algorithms. This section provides a comprehensive overview of these developments, categorized into the contributions and innovations made in three primary areas: the overall domain and taxonomy, the role of agents, and recent significant papers. The classification of existing work in LLM agents for Recommendation and Search is shown in Figure 2.
@@ 4.1.1. Taxonomy Introduction
Deploying LLM agents in search and recommendation domains has given rise to new frameworks and methods. We redefine the taxonomy of these domains, identifying the key domains of LLM agents for RSs as user interaction, representation optimization, system integration, and environment simulation.
@@
Representation Rec4Agentverse (Zhang et al., 2024a) Controls the collaboration between the Intelligent Agent items and the Agent Recommenders.
Representation KGLA (Zeng et al., 2024) Improves user agent memory

### v2

@@ Abstract.
Information technology has profoundly altered the way humans interact with information. The vast amount of content created, shared, and disseminated online has made it increasingly difficult to access relevant information. Over the past two decades, recommender systems and search (collectively referred to as information retrieval systems) have evolved significantly to address these challenges. Recent advances in large language models (LLMs) have demonstrated capabilities that surpass human performance in various language-related tasks and exhibit general understanding, reasoning, and decision-making abilities. This paper explores the transformative potential of LLM agents in enhancing recommender and search systems. We discuss the motivations and roles of LLM agents, and establish a classification framework to elaborate on the existing research. We highlight the immense potential of LLM agents in addressing current challenges in recommendation and search, providing insights into future research directions. This paper is the first to systematically review and classify the research on LLM agents in these domains, offering a novel perspective on leveraging this advanced AI technology for information retrieval. To help understand the existing works, we list the existing papers on LLM agent based recommendation and search at this link: https://github.com/tsinghua-fib-lab/LLM-Agent-for-Recommendation-and-Search.
@@ 1. Introduction
Information technology has fundamentally transformed human life, particularly in how we interact with information. The development of the information age has led to the creation, sharing, and dissemination of vast amounts of content on the internet, making it increasingly challenging for individuals to access relevant information. Against this backdrop, the past two decades have witnessed the evolution of information systems, including Recommender Systems (RSs) (Zhang et al., 2019; Gao et al., 2023b) and search systems (Brin and Page, 1998; Speretta and Gauch, 2005). Typically, search systems passively retrieve relevant content from vast amounts of information with users’ query words, while recommender systems actively generate a list of candidates that users might be interested in based on collected user behavioral data or profile information (Schütze et al., 2008; Aggarwal et al., 2016; Levene, 2011). Some researchers consider recommendation as a specialized form of search, where the understanding of potential user needs is modeled as search terms (Xu et al., 2018).
@@
Recently, large language models (LLMs) have achieved remarkable success and are considered one of the most viable paths toward general artificial intelligence (Achiam et al., 2023). LLMs have demonstrated capabilities similar to or even surpass human performance, not just in language-related tasks, exhibiting general understanding, reasoning, and decision-making abilities (Zhao et al., 2023b). Researchers have leveraged the text processing and commonsense reasoning capabilities of large models to conduct a series of recommendation and search tasks (White, 2024; Zhu et al., 2023; Ai et al., 2023). Furthermore, to overcome the limitations of LLMs, researchers build various LLM agents (Xi et al., 2023). Specifically, LLM agents use LLM as the core and are equipped with functions such as memory management, workflow, input-output interfaces, external tool invocation, and large-small model collaboration, making their intelligence levels closer to human and capable of handling more complex tasks (Zhao et al., 2024a; Park et al., 2023; Zhou et al., 2023c).
@@
For the crucial and fundamental tasks of recommendation and search, researchers increasingly recognize the high value of LLM agents. In this paper, we make the first attempt to provide a systematic and comprehensive review of the research efforts on recommendation and search with LLM agents. We first introduce relevant background knowledge and further discuss the motivations for utilizing LLM agents, as well as how LLM agents address those critical challenges faced by recommendation and search. Next, we establish a taxonomy, categorizing these recent advances according to the role of LLM agents in the overall recommender or search system. We then elaborate on these works and analyze how they construct and use LLM agents. It is worth mentioning there have already been several works discussing the application of LLMs in the fields of IR, which either focus solely on LLMs (Chen et al., 2024a; Bao et al., 2024; Wu et al., 2024b; Gao et al., 2024) or a specific IR application (Peng et al., 2025), but this survey paper significantly differs by focusing on the application of LLM agents across the entire spectrum of IR functionalities.
@@
  • First, this paper is the first to review and organize the research on LLM agents in the research of search and recommendation, which is a new yet fast-growing research field.
@@
The structure of this paper is as follows: In Section 2, we introduce the background knowledge of recommender systems, search systems, LLMs, and LLM agents. In Section 3, we analyze why LLM agents are necessary for recommendation and search. In Section 4, we first establish a coherent and comprehensive taxonomy system and then elaborate on the existing work. In Section 5, we extensively discuss the advancements of embodied LLM agents within the cyber environment, highlighting their potential to inspire the next generation of recommendation and search. In Section 6, we systematically analyze the current problems and significant future directions for recommendation and search based on LLM agents. Finally, we conclude the paper in Section 7.
@@ 2.1.1. Interaction Interface
The rapid development of smart devices and the mobile Internet has exposed users to numerous applications employing recommender systems and search engines. Interacting with a traditional recommender system is simple, where users are mostly unconsciously browsing contents (e.g., on an e-commerce website, or a video sharing platform) provided by the recommender system (Resnick and Varian, 1997). On the other hand, the interaction interface between users and search engines usually consists of a keyword query submission from users (e.g., on a search engine website or a project management software) and a one-round result retrieval from the search engines (Baeza-Yates et al., 1999). Conversational recommender systems and search engines have recently gained attention (Zhang et al., 2018). In this interface, users interact through a chat interface using natural language, iteratively providing additional information or feedback based on previous results, forming a multi-turn dialogue. This conversational interface is more closely aligned with human communication methods and achieves more precise and satisfying results.
@@ 2.1.2. User/Query Modeling
For recommender systems and search engines to retrieve satisfying results, they must understand the users’ intentions. For example, identifying the products a user wants to buy on an e-commerce app or determining relevant websites based on a keyword query. This requires a precise modeling of user characteristics and input queries. In recommender systems, this includes utilizing the user profile (gender, age, preferences, etc.), browsing/clicking/purchasing history, and more (Lops et al., 2011). In search engines, this involves text segmentation, named entity recognition, part-of-speech tagging, relation extraction, etc (Croft et al., 2010). Advanced systems like multi-modal and conversational search engines also incorporate image recognition, multi-turn dialogue comprehension, and other algorithms (Jannach et al., 2021; Wang et al., 2016).
@@ 2.1.3. Item Modeling
Apart from user/query modeling, item modeling also plays an essential role in efficient and extensive retrieval. It gathers information of the items, and preprocesses them into some representation (e.g., a feature vector) for later usage. In recommender systems, item modeling considers item descriptions, attributes, images, buyers, and other related information (Lops et al., 2011; Liu et al., 2024). In traditional search engines, this process is known as Indexing (Croft et al., 2010), where significant keywords or tags from the content of web pages are identified and stored beforehand for faster retrieval. Advanced search engines may also use auxiliary knowledge bases specific to certain domains.
@@ 2.2.1. Large Language Models
Language models play a pivotal role in Natural Language Processing (NLP), primarily used for assessing sentence probabilities and generating grammatically correct text. In its early stages, the development of language models relied on statistical and probability theories such as n-grams, maximum entropy, and Hidden Markov Models, but was limited by its ability to handle long-range dependencies, known as the “curse of dimensionality.” With the rise of deep learning technologies, neural network language models gradually became mainstream due to their robust generalization capabilities and flexible architectural designs. Models like Recurrent Neural Networks (RNNs) (Sherstinsky, 2020), Long Short-Term Memory networks (LSTMs) (Graves and Graves, 2012), and Gated Recurrent Units (GRUs) (Cho et al., 2014) effectively mitigated the challenges of processing long sequences by utilizing internal memory units, thereby better capturing contextual dependencies within sequences and significantly enhancing language model performance.
The introduction of the Transformer model (Vaswani et al., 2017) revolutionized the NLP field by replacing recurrent structures with self-attention mechanisms, thereby improving training speed and parallelism. Concurrently inspired by the pre-training and fine-tuning paradigm from computer vision, language models began leveraging large amounts of unlabeled data for pre-training, followed by task-specific fine-tuning, which gave rise to high-performance models such as BERT (Devlin et al., 2018). With the rapid expansion of training data and significant growth in computational resources, LLMs have established absolute dominance in the NLP. This wave not only symbolizes the comprehensive transformation of NLP from focusing on single-task optimization to multimodal perception and generation but also heralds language models reaching new heights in generation, understanding, and transformation. In the latest research, LLMs such as GPT-4 (Achiam et al., 2023), Claude (Anthropic, 2024), etc., not only perform well in tasks across various NLP tasks such as text generation, multilingual translation, and sentiment analysis but also demonstrate impressive performance in tasks traditionally considered to require deep cognitive abilities like solving mathematical problems, logical reasoning, and strategic planning. This comprehensive performance marks the maturity of artificial intelligence technology, steadily advancing in a direction closer to Artificial General Intelligence (AGI).
@@ 2.2.2. LLM Agents
At the beginning of the 21st century, breakthroughs in machine learning, particularly represented by the rise of deep learning, enabled autonomous learning from massive amounts of data rather than relying solely on pre-set programming rules. This marked a shift in agent research from rule-based methods to data-driven approaches. Deep neural networks significantly enhance the capabilities of intelligent agents in areas such as image recognition, speech processing, and natural language understanding, thus ushering in a new era of Artificial Intelligence (AI) agents. Recently, GPT series (OpenAI, 2022; Achiam et al., 2023) and other major language models have made significant progress in NLP, greatly improving the performance of artificial intelligence agents and demonstrating their profound potential.
@@
  • Perception Module: The perception module is equivalent to the agent’s “senses”, ensuring that the agent can obtain the information needed for decision-making on time. Unlike conventional single-channel information processing systems, LLM agents demonstrate superior capability in seamlessly integrating multimodal information sources, including textual, auditory, visual, and even video data, thereby establishing a more robust and comprehensive perceptual framework (Schumann et al., 2024).
  • Control Module: The control module plays the role of a proxy “brain”, analyzing the collected information, evaluating possible action plans, and selecting the best strategy. In LLM agents, an LLM serves as the core of the control module, enhancing the agent’s understanding and generation ability of language contexts and strengthening its situational adaptability, creativity, and long-term memory functions.
  • Action Module: The action module acts as the agent’s “limbs”, executing decisions to alter environmental states or achieve goals. In the LLM agent, this module is not limited to simple execution actions. Still, it is endowed with deeper functions, which can realize complex tasks such as language generation, tool use, and even concrete actions, greatly expanding the application scope and influence of the agent.
@@
3. Why LLM agents can be used for recommendation and search
@@ 3.2. LLM Agents can Interact with Environment and Integrate Information
In the field of information retrieval, another significant advantage of LLM agents is their ability to interact with environments and integrate results (He et al., 2025; Zhou et al., 2023a). IR often involves multiple data sources and complex retrieval scenarios where traditional retrieval systems may struggle to collect information effectively from different sources. LLM agents with action modules can use various tools, browse web pages, operate mobile apps, and even independently search the internet to find needed information (Schick et al., 2024; Qin et al., 2023). Additionally, LLM agents can go a step further by filtering and summarizing the collected content, presenting users with a clear and concise report instead of delivering raw, unprocessed results like traditional IR systems. For example, when conducting academic literature searches, an LLM agent can interact with multiple academic databases to obtain research outcomes from different fields. It then analyzes and consolidates these results to provide users with a comprehensive literature review. Furthermore, LLM agents can interact with users, continuously adjusting retrieval strategies based on user feedback, thereby enhancing the quality of the retrieval results.
@@ 3.3. LLM Agents Can Serve as User Simulators to Generate Feedback to Information System
IR systems rely on real user feedback for improvement, but conducting experiments on actual users can be costly and negatively impact user experience. The various modules within LLM agents enable them to simulate human perception, memory, and actions, making them well-suited for collecting feedback in place of real users. LLM agents can simulate real user responses when using a product and identifying potential improvements (Zhang et al., 2024g). They can also be used to explore the evolution of users’ mindsets and interests, aiding researchers in better understanding user behavior (Park et al., 2023; Takata et al., 2024). LLM agents can simulate users’ preferences and interact with recommender and search systems. These generated interaction data can be used to train more robust IR models (Sekulić et al., 2024b; Zhang et al., 2024f).
In summary, LLM agents bring several critical capabilities to the table that make them highly suitable for recommendation and search tasks. Their ability to think and plan, interface with various tools, and simulate user behavior makes a significant difference in the user experience of IR in the new era.
@@ 4. Recent Advances of LLM Agents for Recommendation and Search
Recently, significant advancements have been made in the research on combining LLM agents with traditional recommendation and search algorithms. This section provides a comprehensive overview of these developments, categorized into the contributions and innovations made in three primary areas: the overall domain and taxonomy, the role of agents, and recent significant papers. The classification of existing work in LLM agents for Recommendation and Search is shown in Figure 2.
@@ 4.1.1. Taxonomy Introduction
Deploying LLM agents in recommendation and search domains has given rise to new frameworks and methods. We redefine the taxonomy of these domains, identifying the key domains of LLM agents for RSs as user interaction, representation optimization, system integration, and environment simulation.
@@
Representation Rec4Agentverse (Zhang et al., 2024a) Controls the collaboration between the Intelligent Agent items and the Agent Recommenders
Representation KGLA (Guo et al., 2024) Improves user agent memory